OpenStudy (anonymous):

substitute the given values into the given formular and solve for the unknown variable if neccessary round to one decimal place (use the approximation 3.14 for ^) C=2^r; C=18.8 (CIRCUMFERENCE OF A CIRCLE)

OpenStudy (anonymous):

what do u actually want?

OpenStudy (anonymous):

ANSWER

OpenStudy (anonymous):

the formula of circumference is C=2(pi)r

OpenStudy (anonymous):

OKAY TANKS

OpenStudy (anonymous):

dat wrong answer lol

OpenStudy (anonymous):

C=2π∗r r=C/(2π) r=18.8/(2*3.14)

OpenStudy (anonymous):

18.8=2pi r 9.4=pi r r=2.99

OpenStudy (anonymous):

with rounding to one decimal, r=3.0